Alba Progress (Tuesday, 21. July 2009)

7/21/09
Progress in the installation of the synchrotron. Most of the storage ring is now mechanically installed. From our side we have finished the structured cable in the service area and most of the cables for the booster. The cable trays of the storage ring are progressing well too. The Linac area has been partly recabled to its final configuration. There are now more experimental hutches and more progress everywhere. By the way - the quality of the pictures is bad because Fred (Name changed by the editor) has touched the lens and it was super dirty.
The
The "old" beamlines
Roof removed
Roof removed
Installed network cables in the service area
Installed network cables in the service area
Not yet installed cables in the Linac area
Not yet installed cables in the Linac area
Inside the Linac
Inside the Linac
Extra shielding inside the Linac
Extra shielding inside the Linac
Still a little messy
Still a little messy
Vacuum control powered up
Vacuum control powered up
Eric resting his head on a booster bending magnet
Eric resting his head on a booster bending magnet
David and Eric rescrew the connectors for the power and but a nut on the other side of the plate
David and Eric rescrew the connectors for the power and but a nut on the other side of the plate
Watercooling and electical cable in one
Watercooling and electical cable in one
Standard position of the EPS box on the cable tray
Standard position of the EPS box on the cable tray
Nice booster cables and a special cable tray for the EPS cable. This will make the installation of the EPS very neat
Nice booster cables and a special cable tray for the EPS cable. This will make the installation of the EPS very neat
Injection
Injection